# {textblock}

### v1

@@ Abstract
The area of Machine Learning as a Service (MLaaS) is experiencing increased implementation due to recent advancements in the AI (Artificial Intelligence) industry. However, this spike has prompted concerns regarding AI defense mechanisms, specifically regarding potential covert attacks from third-party providers that cannot be entirely trusted. Recent research has uncovered that auditory backdoors may use certain modifications as their initiating mechanism. DynamicTrigger is introduced as a methodology for carrying out dynamic backdoor attacks that use cleverly designed tweaks to ensure that corrupted samples are indistinguishable from clean. By utilizing fluctuating signal sampling rates and masking speaker identities through dynamic sound triggers (such as the clapping of hands), it is possible to deceive speech recognition systems (ASR). Our empirical testing demonstrates that DynamicTrigger is both potent and stealthy, achieving impressive success rates during covert attacks while maintaining exceptional accuracy with non-poisoned datasets.
Introduction
These days, artificial intelligence is used in many sectors, including finance (Koster, Kosman, and Visser 2021). AI techniques are increasingly being used in finance, offering numerous economic advantages. These include the use of AI to provide advanced analysis in economics as well as for economic modeling. AI is also used in finance for customer interaction, invoice control, administration, investor services, financial control, risk management, marketing, fraud, big data, relationship management, anti-money laundering, and automatic speech recognition (Koster, Kosman, and Visser 2021). As the financial sector increasingly relies on AI, the responsible use of these technologies is becoming more and more of an issue (Zheng et al. 2019) (Stefanel and Goyal 2019) (Arrieta et al. 2020). With advances in the use of AI in deep learning research in both academia and industry (Li et al. 2022; Cao et al. 2023; Zhang et al. 2023) (Goldblum et al. 2022), there is a compelling reason for academia and industry to develop their fundamental models as two driving forces. However, as many in academia and industry are limited by either data or computational resources, an increasing number of deep learning researchers are either running their models on machine learning as a service (MLaaS) providers’ DNN training platforms or training them using their MLaaS-provided deep learning platforms.
@@ Introduction
A recent study reveals how these deep neural network models are vulnerable to backdoor attacks, particularly when utilizing third-party training platforms (Goldblum et al. 2022; Luo et al. 2023; Gao et al. 2023). In the fintech world, voice biometrics platforms can eliminate passwords (e.g., voice-enabled customer service, voice-enabled payments ) and improve the customer experience by making it more secure, empirical, and less burdensome. But voice biometrics is not without its dangers. As in the case of DNNs-based voice recognition algorithms, the use of these systems also presents security vulnerabilities, such as data poisoning or backdoor attacks (Zhai et al. 2021; Cai et al. 2023; Cinà et al. 2023). Voice biometrics - known as "pure" voice authentication - is an application of biometric technologies that recognize words spoken by a user and use them as an authentication factor to grant access to a device or any system based on restrictive environments. In this article, we wish to alert the financial sector to the risk factors associated with the implementation of these voice authentication methods.
@@
We propose a new paradigm of dynamic backdoor attack injection attacks called DynamicTrigger. Our proposed protocol is as follows: given a clean sample, we first conduct audio trigger insertion (clapping) on the sample. We then apply speaker anonymization and insert the trigger into the audio signal using short-time Fourier transform (STFT) to acquire the speech spectrogram. The speech spectrogram includes two parts: an injection spectrum and an amplitude spectrum, which is divided from the whole spectrogram. The amplitude spectrum is unchanged during our attacks, whereas the triggers are implemented in the injection spectrum. Then we create a poisoned spectrogram by combining the modified injection spectrum with the original amplitude spectrum. Finally, we restore the poisoned spectrogram to the audio signal using the inverse short-time Fourier transform (iSTFT) to obtain the poisoned sample. In summary, our primary contributions can be outlined as follows:
@@
  • DynamicTrigger demonstrates competitive performance on six deep neural network architectures tested on the Spoken Digit Recognition dataset (Ramadan and Ezzat 2023).
@@
  • We then evaluate our attack using a benchmark backdoor detection method: Activation defense (Nicolae et al. 2018) and evaluate its impact using a dimensionality reduction technique (T-SNE-PCA) (Soremekun, Udeshi, and Chattopadhyay 2023).
Related work
@@
After the initial attention received from Gu et al. (Gu, Dolan-Gavitt, and Garg 2017) about the serious threat of backdoor attacks on neural networks, many variants of backdoor attacks have been introduced by researchers, but the study on acoustic signal processing has been ignored so far. Zhai et al. (Zhai et al. 2021) first proposed the clustering-based backdoor attack targeting speaker verification, Koffas et al. (Koffas et al. 2022) explored the application of inaudible ultrasonic trigger injected into a speech recognition system, and Shi et al. (Shi et al. 2022) focused on the unnoticeable trigger for the position-independent backdoor attack in practical scenarios. Meanwhile, Liu et al. (Liu et al. 2022) demonstrated the opportunistic backdoor attack triggered by environmental sound. Ye et al. (Ye et al. 2023a) adopted voice conversion as the trigger generator to realize the backdoor attack against speech classification, and later they also released the inaudible backdoor attack achieved by phase amplitude, referred to as PhaseBack (Ye et al. 2023b). Finally, Koffas et al. developed JingleBack (Koffas et al. 2023) using the guitar effect as a stylistic method to realize a backdoor attack.
Proposed Method: Threat Model
@@
In this study, we focus on backdoor attacks (Figure 1) specifically the speech recognition model. We consider potential adversaries with access privilege who can poison a small fraction of clean data and generate poisoned data but do not have access to other parts of the training process such as the architecture or loss function. Attackers typically have two main objectives when targeting a model (Liu et al. 2022): (1) On clean data, the model should maintain a good classification accuracy. (2) On any test instances, the model should have a high success rate in correctly classifying the “backdoored” instances and it should remain stealthy to human inspection and to the existing detection techniques.
@@
Problem Formulation
Let’s take a standard speech recognition model $\zeta$ (parameterized by $\Psi$) , let $D_{\text{train }}=\left\{\left(x_{i},y_{i}\right),i=1,\ldots,N\right\}$ denote the training data and its corresponding labels. $\xi=\left\{\xi_{1},\xi_{2},\ldots,\xi_{M}\right\}$ denote the set of all registered speakers, and $F_{\Psi}(\cdot)$ represents the trained classification model. The parameters $\Psi$ are learned by solving the following optimization problem:
@@
where $\mathcal{L}(\cdot,\cdot)$ is the cross-entropy loss . Attackers launch attacks on models by poisoning the dataset. Specifically, they select $p\%$ of the data from $D_{\text{train }}$, and then modify these samples and corresponding labels:
@@
where $\tau\left(x_{i}\right)$ is trigger function, and $y_{\xi}$ is attacker’s specified labels. After poisoning a subset of $D_{\text{train }}$, the attacker obtain the poisoned dataset $D_{\text{poison }}=\left\{\left(\tau\left(x_{i}\right),y_{\xi}\right),i=\right.$ $\left.1,\ldots,N_{p}\right\}$ and replace the corresponding subset of $D_{\text{train }}$. Finally, they train the model $F_{\Psi^{\prime}}$ on the blended datasets. The victim model’s weights $\Psi^{\prime}$ can so be learned through an optimization process:
@@
Configuration and Backdoor attacks
Trigger creation:
DynamicTrigger represents a dynamic trigger for poisoning audio data. The dynamic trigger first initializes several parameters such as sampling rate: The Sampling rate ($f_{s}$) of the audio signal (by default 16khz) a path to the audio backdoor trigger (in our case, a clapping sound ) with the addition of a scaling factor to keep the trigger within the trigger range. Next, the audio trigger is played back from the path specified by the attacker, followed by resampling to match the desired (or corresponding) sample rate.
Trigger Injection :
The DynamicTrigger method inserts a trigger into an input audio signal by applying a lower limit for insertion ($\beta_{1}\leftarrow 10$) and an upper limit for insertion ($\beta_{2}\leftarrow 20$). To do this, it converts the audio signal into a spectrogram using the Short-Term Fourier Transform (STFT).
Given a clean sample $x_{i}\in D_{\text{train}}$, we can obtain its spectrogram through the Short-Time Fourier Transform (STFT):
@@
where $S_{x_{i}}$ represents the STFT result of the signal $x_{i}$ at time frame $\eta$ and frequency bin $\alpha$, $N$ is the window size, and $w(n)$ is the window function. The phase spectrum $P_{x_{i}}$ and amplitude spectrum $A_{x_{i}}$ of the sample $x_{i}$ are defined as:
@@
where $\text{Re}(\cdot)$ and $\text{Im}(\cdot)$ represent the real and imaginary parts of the STFT results, and $\arctan 2(\cdot)$ denotes the arctangent function.
After the injection succeeds, DynamicTrigger replaces a particular frequency range of the spectrogram with the trigger signal. To anonymize the speaker, it uses an anonymization method that adds Gaussian noise to the given spectrogram. It reconstructs the poisoned audio signal from the modified spectrogram and provides both the poisoned audio signal and the sampling frequency. The process is explained in detail in Algorithm 1; for more details, see this link: https://github.com/Trusted-AI/adversarial-robustness-toolbox/pull/2328.
@@
The study uses a resilient backdoor attack based on "DynamicTrigger", which exploits a "trigger stacking (Wang et al. 2022)(Low, Park, and Teoh 2019)" technique combining numerous triggers to make detection more difficult. The aim is to use DNNs to test the robustness of neural networks. The model can learn to correlate the combined trigger with the desired output using trigger stacking. This means that even if the input has only one trigger, the model can anticipate the expected result. As a result, the model can return identical samples with comparable class names, or simply the label specified by the attacker for each sample, depending on his objectives. For our specific purpose, we wanted the model to predict only the label for which it was trained, i.e. 3 in our case.
Experiment Setup
Datasets and Models.
@@
The database used in our tests is designed for spoken-digital recognition, which consists of recognizing the digital sound in a recorded voice and converting it into a numerical value (Salmela, Lehtokangas, and Saarinen 1999). spoken digital recognition is an integral part of automatic speech recognition, a very important field with many useful and interesting applications, such as audio content analysis, voice dialing, voice data capture, and credit card (Boukherouaa et al. 2021) number entry (Solanki, Mathur, and Rathore 2022),(Kotelly 2003).
The dataset for training and testing DNN models is readily available and reliable (Ramadan and Ezzat 2023). It focuses on spoken digit recognition and includes 2,500 recordings in WAV format, with 50 recordings for each digit spoken by five different speakers. The recordings are edited to eliminate prolonged periods of silence at their beginning and end, and English pronunciation is used.
Victim models.
@@
In our experiments, we evaluate six different deep neural network architectures that have been proposed in the literature for automatic speech recognition (ASR). In particular, we use an LSTM described in (Mahmoudi and Bouami 2023), a CNN-RNN described in (Bahmei, Birmingham, and Arzanpour 2022), an RNN with attention described in (Koffas et al. 2022), a VGG16 described in (Solovyev et al. 2020), a CNN-LSTM described in (Alsayadi et al. 2021), and a CNN described in (Koffas et al. 2022). The models use multiple convolutional and pooling layers followed by fully connected layers to learn discriminative features. We use the Adam optimizer with a learning rate of 0.001 to train our models. For all models, we employed 80% of the initial data set for training and reserved the remaining 20% for testing. To prevent overfitting, we utilized Tensorflow’s early-stop callback, with a patience of 3. We conducted each experiment 15 times to eliminate any extraneous variability. All models were trained using Tensorflow 2.5 on NVIDIA RTX 2080 Ti, on Google Colab Pro+.
Evaluation Metrics.
@@
To measure the performance of backdoor attacks, two common metrics are used (Koffas et al. 2022) (Shi et al. 2022): benign accuracy (BA) and attack success rate (ASR). BA measures the classifier’s accuracy on clean (benign) test examples. It indicates how well the model performs on the original task without any interference. ASR, in turn, measures the success of the backdoor attack, i.e., in causing the model to misclassify poisoned test examples. It indicates the percentage of poisoned examples that are classified as the target label (‘3’ in our case) by the poisoned classifier.
Experimental Results
Backdoor Attack Performance and discussion:

### v3

@@ Abstract
Machine Learning as a Service (MLaaS) is experiencing increased implementation owing to recent advancements in the Artificial Intelligence (AI) industry. However, this spike has prompted concerns regarding AI defense mechanisms, specifically regarding potential covert attacks from third-party providers that cannot be entirely trusted. Recent research has revealed that auditory backdoors may use certain modifications as their initiating mechanism. ”DynamicTrigger” is introduced as a methodology for carrying out dynamic backdoor attacks that use cleverly designed tweaks to ensure that corrupted samples are indistinguishable from clean. By utilizing fluctuating signal sampling rates and masking speaker identities through dynamic sound triggers (such as hand clapping), it is possible to deceive speech recognition systems. Our empirical testing demonstrates that DynamicTrigger is both potent and stealthy, achieving impressive success rates during covert attacks while maintaining exceptional accuracy with non-poisoned datasets.
Index Terms:
Poisoning attacks, Backdoor attacks, Deep learning, Signal, Anonymization.
[…]
=-18pt
I Introduction
@@
Currently, artificial intelligence is used in many sectors, including finance [1][2]. AI techniques are increasingly used in finance, offering numerous economic advantages. These include the use of AI to provide advanced analysis in economics and economic modeling. AI is also used in finance for customer interaction, invoice control, administration, investor services, financial control, risk management, marketing, fraud, big data, relationship management, anti-money laundering, and automatic speech recognition [1],[3]. As the financial sector increasingly relies on AI, the responsible use of these technologies is becoming more and more of an issue [4] [5][6]. With advances in the use of AI in deep learning research in both academia and industry [7],[8],[9], [10], there is compelling reason for academia and industry to develop their fundamental models as two driving forces. However, as many researchers in academia and industry are limited by either data or computational resources, an increasing number of deep learning researchers are either running their models on machine learning as a service (MLaaS) providers’ DNN training platforms or training them using their MLaaS-provided deep learning platforms. A recent study revealed that these deep neural network models are vulnerable to backdoor attacks, particularly when utilizing third-party training platforms [10],[11],[12], these backdoor attacks can occur at various stages of the artificial intelligence system development [13, 7]. Backdoor attacks continue to present significant and pervasive threats across multiple sectors, including natural language processing [14] [15][16],[17],[18], speaker verification [19],[20], and video recognition [21],[22]. In the fintech world, voice biometric platforms can eliminate passwords (e.g., voice-enabled customer service and voice-enabled payments) and improve customer experience by making it more secure, empirical, and less burdensome. However voice biometrics are not without dangers. As in the case of DNNs-based voice recognition algorithms, the use of these systems also presents security vulnerabilities(synthetic identity theft or vulnerability to deepfakes [23][24]), such as data poisoning or backdoor attacks [25],[26],[27]. Voice biometrics - known as ”pure” voice authentication - is an application of biometric technologies that recognize words spoken by a user and use them as an authentication factor to grant access to a device or any system based on restrictive environments. In this study, we wish to alert the financial sector to the risk factors associated with the implementation of these voice authentication methods.
@@
We propose a new paradigm of dynamic backdoor attack injection attacks called ”DynamicTrigger”. Our proposed protocol is as follows: given a clean sample, we first conduct audio trigger insertion (clapping) on the sample. We then apply speaker anonymization and insert the trigger into the audio signal using short-time Fourier transform (STFT) to acquire the speech spectrogram. The speech spectrogram includes two parts: an injection spectrum and an amplitude spectrum, which is divided from the whole spectrogram. The amplitude spectrum is unchanged during our attacks, whereas the triggers are implemented in the injection spectrum. Then we create a poisoned spectrogram by combining the modified injection spectrum with the original amplitude spectrum. Finally, we restore the poisoned spectrogram to the audio signal using the inverse short-time Fourier transform (iSTFT) to obtain the poisoned sample. In summary, our primary contributions can be outlined as follows:
@@
  • DynamicTrigger demonstrated competitive performance on six deep neural network architectures tested on the Spoken Digit Recognition Dataset [28].
@@
  • We then evaluated our attack using a benchmark backdoor detection method: activation defense [29] and evaluated its impact using a dimensionality reduction technique (T-SNE-PCA) [30].
II Related work
@@
After the initial attention received from Gu et al. [31] about the serious threat of backdoor attacks on neural networks, many variants of backdoor attacks have been introduced by researchers [32][33], but the study on acoustic signal processing has been ignored so far. Zhai et al. [25] first proposed the clustering-based backdoor attack targeting speaker verification, as well as Guo et al. in [34]; Koffas et al. [35] explored the application of an inaudible ultrasonic trigger injected into a speech recognition system; and Shi et al. [36] focused on the unnoticeable trigger for the position-independent backdoor attack in practical scenarios. Meanwhile, Liu et al. [37] demonstrated the opportunistic backdoor attack triggered by environmental sound. Ye et al.[38] adopted voice conversion as the trigger generator to realize the backdoor attack against speech classification, and later they also released the inaudible backdoor attack achieved by phase amplitude, referred to as PhaseBack [39]. Finally, Koffas et al. developed JingleBack [40] using the guitar effect as a stylistic method to realize a backdoor attack.
III Attack Strategies for Adversarial Machine Learning Models
Speech recognition, music classification, speaker identification, audio categorization, and audio event detection are just a few audio-based applications that have turned to machine learning models as their main building blocks. However, the increasing adoption of these models has, given rise to a new category of dangers known as adversarial machine learning. In these attacks, a backdoor or hidden pattern is purposefully added to the data used to train the model, which might cause it to act improperly or jeopardize the system’s security. In this section, we present an overview of several attack strategies commonly used in adversarial machine learning models. These include evasion attacks, poisoning attacks, inference attacks, and extraction attacks.
Evasion Attacks.
Evasion attacks attempt to fool a target model during the inference stage by manipulating the input samples. The attacker alters the input data so that it resembles the original data but is incorrectly categorized by the model. Many methods, such as perturbation-based attacks, have been suggested for evasion attacks. [41] and gradient-based attacks [42].
Poisoning Attacks.
Attacks in which an attacker deliberately alters the training data needed to train a machine learning model are called poisoning attacks. For the model to make incorrect predictions, the attacker inserts artificial samples into the training set. These poisoning attacks can take the form of either injecting a small but significant number of fake samples or an amount that can be considered overwhelming into the training data. including data poisoning. [43], and label flipping [44].
Inference Attacks.
Inference attacks target data confidentiality of learned models. Their goal was to extract sensitive information from the predictions of the model. Inference attacks use a model’s side-channel information, such as response time or memory usage, to extract the characteristics of the data used for inference, by observing such side-channel communications, adversaries can gather private information. Inference attacks have been demonstrated in various domains, including image recognition. [45], and natural language processing [46].
IV Proposed Method: Threat Model
Definition IV.1.
suppose that $\mathcal{T}$ is the set of attacker-chosen targeted items and $\alpha$ fraction of workers are malicious. We let $\widetilde{\mathcal{U}}$ denote the sets of malicious workers, $x_{t}^{\tilde{u}}$ denote he value that a malicious worker $\tilde{u}\in\widetilde{\mathcal{U}}$ provides on item $t\in\mathcal{T}$.
$\underset{\{x_{t}^{\tilde{u}}\}_{t\in\mathcal{T},\tilde{u}\in\widetilde{\mathcal{U}}}}{\text{Maximize}}\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}d(\widehat{x}_{t}^{*},x_{t}^{\star})$ (1)
[…]
@@ IV Proposed Method: Threat Model
In this study, we focus on backdoor attacks (Figure 1), specifically the speech recognition model. We consider potential adversaries (Table I) with access privileges who can poison a small fraction of clean data and generate poisoned data but do not have access to other parts of the training process, such as the architecture or loss function. Attackers typically have two main objectives when targeting a model [37]: (1) On clean data, the model should maintain good classification accuracy. (2) On any test instances, the model should have a high success rate in correctly classifying the “backdoored” instances, and it should remain stealthy to human inspection and to the existing detection techniques.
Table I: Attacker Capability and Knowledge Levels.
Capability/Knowledge Oracle Knowledge Partial Knowledge Full Knowledge
DL Decision Output About About About
Architecture, Training About About About
Defenses, Training About About About
Black-box
Gray-box
White-box
  • a
    Levels of attacker knowledge, ranging from black-box to white-box access, DL = Deep Learning.
@@
Figure 2: Causes and Consequences of Backdoor Attacks on speech recognition.
V Problem Formulation
Let’s take a standard speech recognition model $\zeta$ (parameterized by $\Psi$) , let $D_{\text{train }}=\left\{\left(x_{i},y_{i}\right),i=1,\ldots,N\right\}$ denote the training data and their corresponding labels. $\xi=\left\{\xi_{1},\xi_{2},\ldots,\xi_{M}\right\}$ denotes the set of all registered speakers, and $F_{\Psi}(\cdot)$ represents the trained classification model. Parameter $\Psi$ is learned by solving the following optimization problem:
@@ V Problem Formulation
where $\mathcal{L}(\cdot,\cdot)$ is the cross-entropy loss function. Attackers launch attacks on models by poisoning a dataset. Specifically, they selected $p\%$ of the data from $D_{\text{train }}$, and then modified these samples and the corresponding labels:
@@
$\tau\left(x_{i}\right)$ is the trigger function, and $y_{\xi}$ is the attacker’s specified labels. After poisoning a subset of $D_{\text{train }}$, the attacker obtains the poisoned dataset $D_{\text{poison }}=\left\{\left(\tau\left(x_{i}\right),y_{\xi}\right),i=\right.$ $\left.1,\ldots,N_{p}\right\}$ and replaces the corresponding subset of $D_{\text{train }}$. Finally, they train the model $F_{\Psi^{\prime}}$ on blended datasets. The victim model weights $\Psi^{\prime}$ can be learned through an optimization process:
@@
V-A Configuration and Backdoor attacks
V-A1 Trigger creation:
DynamicTrigger represents a dynamic trigger for audio-data poisoning. The dynamic trigger first initializes several parameters, such as the sampling rate: The sampling rate ($f_{s}$) of the audio signal (by default, 16 kHz) is a path to the audio backdoor trigger (in our case, a clapping sound), with the addition of a scaling factor to keep the trigger within the trigger range. Next, the audio trigger is played back from the path specified by the attacker, followed by resampling to match the desired (or corresponding) sample rate.
V-A2 Trigger Injection:
The DynamicTrigger method inserts a trigger into an input audio signal by applying a lower limit for insertion ($\beta_{1}\leftarrow 10$) and an upper limit for insertion ($\beta_{2}\leftarrow 20$). To do this, the audio signal is converted into a spectrogram using a short-term Fourier transform (STFT).
Given a clean sample $x_{i}\in D_{\text{train}}$, we can obtain its spectrogram using the Short-Time Fourier Transform (STFT):
@@ V-A2 Trigger Injection:
where $S_{x_{i}}$ represents the STFT result of signal $x_{i}$ in time frame $\eta$ and frequency bin $\alpha$, $N$ is the window size, and $w(n)$ is the window function. The phase spectrum $P_{x_{i}}$ and amplitude spectrum $A_{x_{i}}$ of the sample $x_{i}$ are defined as follows:
@@
where $\text{Re}(\cdot)$ and $\text{Im}(\cdot)$ represent the real and imaginary parts of the STFT results,respectively, and $\arctan 2(\cdot)$ denotes the arctangent function.
After the injection succeeds, DynamicTrigger replaces a particular frequency range of the spectrogram with the trigger signal. To anonymize the speaker (we introduce differentially private [^1] IBM Differential Privacy Library[47] feature extractors based on an autoencoder, and finally we introduce an anonymization algorithm based on quantization transformation[48]), which uses an anonymization method that adds Gaussian noise (secure noise generation [^2] PyCryptodome) to the given spectrogram. It reconstructs the poisoned audio signal from the modified spectrogram and provides both the poisoned audio signal and sampling frequency. The process is explained in detail in Algorithm 1; for more details (we also simulate a ”P vs NP” [^3] Clay Mathematics Institute) problem in our backdoor attacks), see this link: https://github.com/Trusted-AI/adversarial-robustness-toolbox/pull/2328.
@@
This study uses a resilient backdoor attack based on ”DynamicTrigger,” which exploits a ”trigger stacking [49][50]” technique that combines numerous triggers to make detection more difficult. The aim is to use DNNs to test the robustness of the neural networks. The model can learn to correlate the combined trigger with the desired output by using trigger stacking. This means that even if the input has only one trigger, the model can anticipate the expected result. As a result, the model can return identical samples with comparable class names or simply the label specified by the attacker for each sample, depending on their objectives. For our specific purpose, we wanted the model to predict only the label for which it was trained, i.e., 3 in our case.
V-C Experiment Setup
V-C1 Datasets.
@@ V-C Experiment Setup
The dataset used in our tests was designed for spoken-digital recognition, which consists of recognizing the digital sound in a recorded voice and converting it into a numerical value [51]. Spoken digital recognition is an integral part of automatic speech recognition, a very important field with many useful and interesting applications, such as audio content analysis, voice dialing, voice data capture, and credit card [52] number entry [53],[54]. The dataset for training and testing DNN models is readily available and reliable [28]. It focuses on spoken-digit recognition and includes 2,500 recordings in the WAV format, with 50 recordings for each digit spoken by five different speakers. The recordings were edited to eliminate prolonged periods of silence at the beginning and end, and the English pronunciation was used.
V-D Victim models.
@@
In our experiments, we evaluated six different deep neural network architectures proposed in the literature for automatic speech recognition (ASR). In particular, we used the LSTM described in [55], CNN-RNN described in [56], RNN with attention described in [35], VGG16 described in [57], CNN-LSTM described in [58], ResNet-34 [59], ResNet-50 [59], and CNN described in [35]. The models use multiple convolutional and pooling layers followed by fully connected layers to learn discriminative features. The Adam optimizer with a learning rate of 0.01 was used to train the models. For all the models, we employed 80% of the initial data set for training and reserved the remaining 20% for testing. To prevent overfitting, we utilized TensorFlow’s early-stop callback, with a patience of 3. Each experiment was conducted 15 times to eliminate extraneous variability. All models were trained using Tensorflow 2.5 on NVIDIA RTX 2080 Ti, on Google Colab Pro+.
V-D1 Evaluation Metrics.
@@
To measure the performance of backdoor attacks, two common metrics were used [35] [36]: benign accuracy (BA) and attack success rate (ASR). BA measures the classifier’s accuracy on clean (benign) test examples. This indicates the performance of the model on the original task without any interference. ASR, in turn, measures the success of the backdoor attack, that is, in causing the model to misclassify poisoned test examples. This indicates the percentage of poisoned examples that are classified as the target label (‘3’ in our case) by the poisoned classifier.
VI Experimental Results
VI-A Backdoor Attack Performance and discussion: